### v1

# DynaSolidGeo: A Dynamic Benchmark for Genuine Spatial Mathematical Reasoning of VLMs in Solid Geometry

Abstract
Solid geometry problem solving demands spatial mathematical reasoning that integrates spatial intelligence and symbolic reasoning. However, most existing multimodal mathematical reasoning benchmarks focus primarily on 2D plane geometry, rely on static datasets prone to data contamination and memorization, and evaluate models solely by final answers, overlooking the reasoning process. To address these limitations, we introduce DynaSolidGeo, the first dynamic benchmark for evaluating genuine spatial reasoning in Vision-Language Models (VLMs). Constructed through a semi-automatic annotation pipeline, DynaSolidGeo contains 503 expert-curated seed questions that can, in principle, dynamically generate an unbounded number of diverse multimodal text-visual instances. Beyond answer accuracy, we incorporate process evaluation based on expert-annotated reasoning chains to measure logical validity and causal coherence. Experiments across representative open-source and closed-source VLMs reveal large performance gaps, severe degradation in dynamic settings, and poor performance on tasks requiring high-level spatial intelligence, such as mental rotation and visualization. The code and dataset are available at DynaSolidGeo.
| Benchmarks | Language | S.G. Size (Proportion) | S.G. Category | Level | Static/Dynamic | Modality | Metric |
| GeoQA [8] | EN&CN | 0 (0.0%) | - | K. | Static | Text&Image | A.E. |
| PGPS9K [42] | EN | 0 (0.0%) | - | K. | Static | Text&Image | A.E. |
| MMMU-MATH [40] | EN | 0 (0.0%) | - | U. | Static | Text&Image | A.E. |
| GeoEval [41] | EN | 100 (2.0%) | - | K. | Static | Text&Image | A.E. |
| MATH-Vision [32] | EN | 263 (8.7%) | - | K.&U. | Static | Text&Image | A.E. |
| OlympiadBench [15] | EN&CN | 784 (9.2%) | - | C. | Static | Text&Image | A.E. |
| MathVerse [43] | EN | 119 (15.1%) | - | K. | Static | Text&Image | A.E. |
| GeoSense [38] | EN&CN | 350 (20.0%) | 2 | K. | Static | Text&Image | A.E.&P.E. |
| SolidGeo [33] | EN&CN | 3113 (100.0%) | 8 | K.&U.&C. | Static | Text&Image | A.E. |
| GeoLaux [13] | CN | 0 (0.0%) | - | K. | Static | Text&Image | A.E.&P.E. |
| DynaMath [47] | EN&CN | 15 seed questions (3.0%) | - | K.&U. | Dynamic | Text&Image | A.E. |
| DynaSolidGeo (Ours) | EN&CN | 503 seed questions (100.0%) | 8 | K.&C. | Dynamic | Text&Image&Video | A.E.&P.E. |
1 Introduction
Geometry problem solving has long played a central role in mathematical reasoning, requiring integrating visual understanding and symbolic reasoning across complex graphic and textual contexts [38]. According to structural properties, geometry can be categorized into plane geometry and solid geometry. Compared to plane geometry, solid geometry imposes substantially higher spatial mathematical reasoning ability, as reasoning in three dimensions entails spatial intelligence, including spatial perception, spatial relation, spatial orientation, spatial rotation, and spatial visualization that goes beyond two-dimensional recognition [14, 29, 16, 22]. Such tasks remain difficult even for human learners [7], and represent a formidable open challenge for current AI systems.
In parallel, recent years have witnessed remarkable progress in multimodal large language models (MLLMs). Building on the successes of foundation models, vision-language models (VLMs) [23, 6, 11, 27, 5, 4, 17] have rapidly advanced the state of the art in a wide spectrum of multimodal understanding tasks. Among these tasks, multimodal mathematical reasoning has emerged as a challenging yet vibrant frontier, with benchmarks such as GeoQA [8], MathVista [25], and GeoSense [38] exposing both the promise and the limitations of current VLMs. These carefully designed benchmarks have played a pivotal role in advancing the field, providing standardized evaluation and catalyzing iterative improvements in model design and training paradigm.
Despite this progress, current multimodal mathematical reasoning benchmarks exhibit critical limitations. First, the vast majority of existing geometry-related benchmarks focus on plane geometry or diagram-based word problems, leaving solid geometry, which places heightened demands for spatial intelligence, largely underexplored. For example, PGPS9K [42] contains more than 9,000 plane geometry questions but no solid geometry items, and in GeoEval [41], tasks involving solid geometry constitute merely 2% of the benchmark. Second, nearly all existing multimodal mathematical reasoning benchmarks are static, relying on fixed and finite test sets that are susceptible to data contamination and memorization. Recent analyses demonstrate that large models can memorize and regurgitate benchmark data [26, 12, 21, 28, 10], and some studies show that decontaminated re-releases often lead to substantial drops in performance [44], confirming that static evaluation may significantly overestimate true reasoning and generalization [31]. Similar concerns have motivated dynamic evaluations in coding [18, 45] and general-purpose QA [34, 35], indicating a community-wide shift towards dynamic, contamination-resistant evaluation protocols [9]. Moreover, most existing multimodal mathematical reasoning benchmarks [25, 33, 32, 41, 40, 42, 47] evaluate models solely based on answer accuracy, which allows models suffering from data contamination or over-reliance on memorization to appear strong, while failing to reveal their genuine reasoning ability.
To address these limitations, we present DynaSolidGeo, a new benchmark for the dynamic evaluation of VLMs’ genuine spatial mathematical reasoning in solid geometry. Unlike existing static resources, DynaSolidGeo consists of 503 seed questions of solid geometry problem solving, each represented by a Python program paired with a corresponding MATLAB program. With the correctness of the question guaranteed, each seed question is parameterized: textual variables in the question statement (e.g., endpoint labels, side lengths, areas, volumes, ratio) as well as rendering parameters of the solid geometry (e.g., camera viewpoints) can all be randomized. By supplying different random seeds, DynaSolidGeo can in principle generate an unbounded number of question-answer instances, where each instance can optionally include two visual versions: a randomized-view image and a 360-degree rotation video. The seed questions of DynaSolidGeo are drawn from diverse and authoritative sources, including China’s Gaokao examinations, international mathematics competitions, and widely used training materials for competition preparation. Together, they cover nearly all major categories of high-school and competition-level solid geometry problems (eight in total), including positional relations, angle, length, area, and volume calculations, as well as counting, dynamic, and folding tasks. Moreover, we move beyond answer-only evaluation by incorporating process-level assessment grounded in expert-annotated reasoning chains. Through Answer Accuracy (AA), Process Score (PS), and Process-Qualified Accuracy (PA), we jointly measure answer correctness, reasoning quality, and reasoning-qualified accuracy, offering a more faithful reflection of VLMs’ genuine spatial mathematical reasoning ability. To ensure reliability, all solutions are expert-annotated by undergraduates and graduate students from the School of Mathematical Sciences, Peking University, including Chinese Mathematical Olympiad (CMO) gold medalists. A comparative summary with related benchmarks is provided in Table 1.
We evaluate a range of mainstream, latest closed- and open-source VLMs on DynaSolidGeo. Experiments reveal a clear gap between most open-source and closed-source VLMs. Notably, nearly all models struggle with Counting problems, highlighting the lack of higher-order spatial intelligence, such as mental rotation and spatial visualization. Compared to the static source-question dataset, models exhibit a significant performance drop on DynaSolidGeo (up to 20.4% for Claude-Sonnet-4.5), exposing potential data contamination and memorization effects. Furthermore, the additional metric degradation after introducing process evaluation indicates that previous static, answer-only benchmarks likely overestimated model capabilities, whereas DynaSolidGeo provides a more faithful and comprehensive evaluation of genuine spatial mathematical reasoning ability. In summary, our contributions are as follows:
We design a semi-automatic data annotation pipeline for the seed question annotation of solid geometry problems, which minimizes human involvement without compromising annotation correctness or usability.
We propose DynaSolidGeo, the first dynamic benchmark for solid geometry problem solving, consisting of 503 carefully curated seed questions that can, in principle, automatically generate an unbounded number of diverse question instances across multiple geometry categories.
We introduce a process evaluation using expert-annotated reasoning chains that, together with answer evaluation, provides a holistic measure of VLMs’ genuine spatial mathematical reasoning capability.
We evaluate a series of popular and SOTA VLMs on DynaSolidGeo to gain deeper insights into their spatial mathematical reasoning abilities and conduct extensive analyses, including revealing potential data contamination and memorization phenomenon on static datasets.
2 Related Work
2.1 Multimodal Mathematical Reasoning Benchmarks.
Recent years have witnessed the emergence of multimodal benchmarks that evaluate mathematical reasoning in visually grounded settings. Early efforts include TQA [19] and Geometry3K [24] introduced multimodal reasoning tasks involving diagram-based science and geometry word problems with accompanying 2D visuals. More recent benchmarks, such as GeoQA [8], PGPS9K [42], MMMU-MATH [40], GeoEval [41], MATH-Vision [32], OlympiadBench [15], MathVerse [43], GeoSense [38], and GeoLaux [13] have broadened coverage to thousands of multimodal math problems. However, the vast majority of these resources concentrate on plane geometry and 2D diagrammatic reasoning, leaving solid geometry largely underexplored. A few datasets have attempted to move toward 3D: SolidGeo [33] explicitly targets solid geometry but remains static datasets vulnerable to contamination and memorization; DynaMath [47] introduces dynamic instance generation, but solid geometry is barely represented, with only 15 problems (3%) in the dataset. In contrast, DynaSolidGeo fills this gap with scalable and dynamic solid-geometry coverage.
2.2 Vision-Language Models.
Recent vision–language models (VLMs) such as BLIP-2 [20], Flamingo [3], and LLaVA [23] combine pretrained large language models with visual encoders, enabling open-ended multimodal reasoning and instruction following. Building on this paradigm, the latest generation of VLMs has rapidly advanced in scale, architecture, and reasoning capability. The closed-source models include GPT-5 family [27], Gemini-2.5 family [11], and Claude-Sonnet-4.5 [5], which feature deeply integrated multimodal backbones and enhanced reasoning modules. In parallel, the open-source community has introduced competitive alternatives such as LLaVA-OneVision-1.5 family [4], GLM-4.1V-9B-Thinking [17], Llama-4-Maverick-17B-Instruct [2], InternVL3.5-8B [37], DeepSeek-VL2 [36], and the Qwen3-VL family [30], which push the frontier of visual capabilities. Yet their spatial mathematical reasoning ability remains underexplored, motivating our evaluation on DynaSolidGeo.
3 DynaSolidGeo
We propose DynaSolidGeo, a dynamic multimodal benchmark for spatial mathematical reasoning in solid geometry, which consists of 503 expert-annotated seed questions that can expand into unbounded question–answer instances with randomized text, images, and 360-degree rotation videos, by inputting a random seed.
3.1 Data Collection
The seed questions of DynaSolidGeo are drawn from diverse and authoritative sources to ensure both breadth and rigor. Specifically, we collect 503 solid geometry questions (referred to as source questions) from three major categories: 1) China’s Gaokao examinations from 2014 to 2025 (11 years), 2) international mathematics competitions such as the American Invitational Mathematics Examination (AIME), the American Mathematics Competitions (AMC), and the American High School Mathematics Examination (AHSME), and 3) high-level preparation and training materials, including competition handbooks and advanced supplementary textbooks. These sources cover nearly the full spectrum of high-school and competition-level solid geometry categories, encompassing positional relationships, angles, distances, area and volume computation, as well as combinatorial counting, dynamic scenarios, and folding/unfolding problems (see Table 2).
3.2 Data Annotation Pipeline
We design a semi-automatic seed-question annotation pipeline that aims to minimize manual labeling costs while ensuring the correctness and availability of the generated programs. Compared with a fully manual annotation process, this approach substantially reduces human effort. At the same time, in contrast to a fully automatic procedure, it preserves accuracy and reliability in handling complex solid geometry questions. As shown in Figure 1, our data annotation process is divided into the following components:
1.
Expert-Guided Parametrization and Visualization: Here we follows a human-in-the-loop strategy, where human experts collaborate with large models to create a JSON annotation and a MATLAB program for each source question:
JSON-Based Question Parametrization: For each collected source question, mathematics experts parameterize the question statement by converting fixed values into variable parameters (e.g., endpoint labels, side lengths, areas, volumes, ratios) using f-string syntax, while ensuring correctness and availability. Corresponding answer is also expressed in terms of these variable parameters. An example of parameterized variables is highlighted in red in Fig. 1. Additional metadata (e.g., category and difficulty level) is also included and stored in JSON format.
MATLAB-Based Geometry Visualization: MATLAB experts implement programs that render each solid geometry image and video for each source question.
2.
Automated Python Program Synthesis:
Answer Function Generation: With the assistance of the large language model, parametrized answers in JSON are converted into Python functions that dynamically compute results.
Parameterized Program Construction: A rule-based script automatically assembles parameterized Python programs from the parametrized questions and answer functions. By inputting a random seed, the parameterized Python program randomizes both the MATLAB camera parameters (i.e., azimuth and elevation) and variable parameters in the parameterized question, and finally outputs a JSON entry of the instantiated question along with a MATLAB invoke command.
3.
Automated MATLAB Program Synthesis: Each MATLAB visualization program is automatically converted into a parameterized version by a rule-based script, aligned with the annotated JSON specification. These programs can be directly invoked by the MATLAB commands generated in the previous step, enabling dynamic rendering of figures and videos consistent with the instantiated question parameters.
4.
Expert Verification: Final human checks ensure correctness, consistency, and usability of seed questions.
Overall, each seed question is associated with a parameterized Python program for generating the textual description and a parameterized MATLAB program for rendering the corresponding figures and videos. By inputting a random seed, each seed question can be instantiated into a concrete question instance.
3.3 Statistics
Table 2 summarizes the detailed statistics of the DynaSolidGeo dataset. In total, the benchmark contains 503 curated seed questions, all newly constructed for this work. The questions span a diverse set of solid geometry problem categories, including Positional relationship determination (PD, 11.7%), Angle calculation (AN, 20.5%), Length and distance calculation (LC, 13.1%), Area calculation (AR, 11.3%), Volume calculation (VC, 10.3%), Counting problems (CP, 7.0%), Dynamic or moving-point problems (DM, 13.1%), and Folding and unfolding problems (FP, 12.9%). The distribution across difficulty levels is reasonably balanced, with 27.2% easy, 57.7% medium, and 15.1% hard questions. Regarding question types, previous benchmarks (e.g., DynaMath[47], SolidGeo [33], GeoSense[38], MathVerse[43]) generally include multiple-choice questions. Such options inevitably provide strong hints to the models, thereby reducing the difficulty and making it difficult to assess their reasoning ability accurately. In contrast, we rewrite the multiple-choice and proof questions from the source data into fill-in-the-blank formats in our work. As a result, DynaSolidGeo consists of 88.3% numerical questions and 11.7% free-form questions, posing greater challenges to the reasoning ability of VLMs.
Figure 2 shows the distribution of the number of variable parameters contained in the seed questions of DynaSolidGeo. The variable parameters include camera parameters (i.e., azimuth and elevation), endpoint labels, side lengths, areas, volumes, ratios, and so on. As illustrated, the seed questions exhibit substantial variability, highlighting the richness and flexibility of our benchmark design.
| Statistic | Number |
| Total seed questions | 503 |
| - Newly curated questions | 503 (100.0%) |
| Categories | |
| - Positional relationship determination (PD) | 59 (11.7%) |
| - Angle calculation (AN) | 103 (20.5%) |
| - Length and distance calculation (LC) | 66 (13.1%) |
| - Area calculation (AR) | 57 (11.3%) |
| - Volume calculation (VC) | 52 (10.3%) |
| - Counting problems (CP) | 35 (7.0%) |
| - Dynamic or moving-point problems (DM) | 66 (13.1%) |
| - Folding and unfolding problems (FP) | 65 (12.9%) |
| Levels | |
| - Easy | 137 (27.2%) |
| - Medium | 290 (57.7%) |
| - Hard | 76 (15.1%) |
| Question types | |
| - Numerical questions | 444 (88.3%) |
| - Free-form questions | 59 (11.7%) |
3.4 Evaluation Metrics
Unlike most existing studies that only assess the correctness of final answers, we additionally introduce process evaluation to holistically assess the model’s genuine spatial reasoning capacity. DynaSolidGeo consists of seed questions, from which we instantiate batches of question instances by sampling with different random seeds. Building on these instances, we design the following evaluation metrics to systematically evaluate model performance.
3.4.1 Answer Accuracy
We use the Answer Accuracy (AA) to measure the correctness of the model’s answers, i.e., the proportion of final answers that are correct on average:
where and denote the rule-extracted answer and the corresponding ground truth of the -th question in the -th batch, respectively.
3.4.2 Process Score
To more accurately assess the quality of the model’s reasoning process, we annotate not only parameterized question statements and answers, but also parameterized reference reasoning processes. Each reasoning process is evaluated using LLM as a judge according to the following criteria:
Logical Alignment: The reasoning presents a coherent derivation whose steps consistently lead to the stated result, with matching variables/units and without any unjustified conclusion jumps.
No Extraneous Information: The reasoning does not rely on unseen quantities or unsupported facts as essential premises (standard geometric axioms and theorems are allowed).
Use of Key Dependencies: The reasoning explicitly invokes the key geometric relations specified in the problem (e.g., parallelism, similarity, perpendicularity, collinearity, ratios, or angle constraints), rather than skipping these conditions and merely reporting the final result.
Based on these criteria, we introduce the process evaluation metric, termed the Process Score (PS):
In Equation 2, denotes the process score of the -th question in the -th batch scored by the judge model. A higher PS corresponds to a reasoning process that is more accurate, coherent, and of higher quality.
3.4.3 Process-Qualified Accuracy
Although some questions are answered correctly, the reasoning process behind them may not be logically accurate, coherent, or supportive of the final correct answer. As a result, such “hallucination” cases inflate the evaluation of the model’s spatial mathematical reasoning ability. To address this, we propose a new composite metric, Process-Qualified Accuracy (PA), which combines Answer Accuracy and Process Score to more accurately measure the model’s true spatial mathematical reasoning capability:
We believe that a model can only truly possess the ability to solve a problem when its reasoning process is logically accurate, coherent, and supports the final answer, rather than merely relying on the correctness of the final answer alone.
| Model | PD | AN | LC | AR | VC | CP | DM | FP | ALL |
| AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | |
| Closed-sourced MLLMs | |||||||||
| GPT-5-Nano | 39.5 / - / - | 54.0 / - / - | 56.1 / - / - | 71.9 / - / - | 71.2 / - / - | 5.7 / - / - | 53.0 / - / - | 42.6 / - / - | 51.4 / - / - |
| GPT-5 | / - / - | / - / - | / - / - | / - / - | / - / - | 20.0 / - / - | / - / - | / - / - | / - / - |
| Gemini-2.5-Flash | 44.1 / 42.9 / 42.9 | 48.2 / 45.9 / 45.6 | 60.1 / 57.8 / 58.1 | 63.7 / 61.0 / 61.4 | 55.8 / 53.7 / 53.9 | 16.2 / 16.2 / 16.2 | 61.6 / 59.8 / 60.1 | 34.9 / 33.6 / 33.9 | 49.6 / 47.8 / 47.9 |
| Gemini-2.5-Pro | 71.8 / 61.2 / 54.8 | 52.4 / 43.6 / 39.2 | 69.7 / 64.6 / 61.6 | 71.3 / 67.5 / 66.1 | 76.9 / 69.7 / 65.4 | / 30.5 / 30.5 | 63.1 / 60.0 / 57.6 | 56.4 / 50.4 / 47.7 | 62.0 / 55.9 / 52.6 |
| Claude-Sonnet-4.5 | 43.5 / 34.7 / 32.2 | 26.5 / 25.0 / 24.0 | 37.9 / 34.5 / 32.8 | 50.3 / 48.7 / 49.1 | 53.8 / 49.8 / 47.4 | 6.7 / 6.0 / 5.7 | 26.3 / 24.2 / 23.2 | 15.4 / 13.5 / 12.8 | 32.7 / 29.7 / 28.6 |
| Open-sourced VLMs | |||||||||
| LLaVA-OneVision-1.5-4B-Instruct | 11.9 / - / - | 4.5 / - / - | 7.6 / - / - | 15.8 / - / - | 10.3 / - / - | 1.9 / - / - | 4.0 / - / - | 0.0 / - / - | 6.8 / - / - |
| LLaVA-OneVision-1.5-8B-Instruct | 17.5 / - / - | 1.9 / - / - | 7.6 / - / - | 2.9 / - / - | 2.6 / - / - | 1.0 / - / - | 5.6 / - / - | 3.1 / - / - | 5.2 / - / - |
| GLM-4.5V | 49.7 / - / - | 31.4 / - / - | 42.9 / - / - | 57.3 / - / - | 50.6 / - / - | 7.6 / - / - | 48.0 / - / - | 12.8 / - / - | 38.1 / - / - |
| GLM-4.1V-9B-Thinking | 29.9 / 27.3 / 26.6 | 22.7 / 21.7 / 21.7 | 33.3 / 30.7 / 31.3 | 44.4 / 42.7 / 43.3 | 41.0 / 37.8 / 39.1 | 2.9 / 2.9 / 2.9 | 26.8 / 25.5 / 25.3 | 5.6 / 4.6 / 4.1 | 26.2 / 24.6 / 24.7 |
| Llama-3.2-90B-Vision-Instruct | 35.6 / 27.1 / 23.2 | 14.2 / 10.1 / 8.4 | 25.8 / 22.6 / 21.2 | 49.7 / 45.6 / 43.9 | 39.7 / 34.0 / 31.4 | 1.9 / 1.2 / 1.0 | 9.1 / 6.7 / 5.6 | 8.2 / 5.0 / 3.1 | 22.6 / 18.5 / 16.6 |
| Llama-4-Maverick-17B-Instruct | 36.7 / 27.1 / 21.5 | 13.9 / 10.4 / 8.4 | 24.7 / 21.8 / 20.7 | 46.8 / 44.3 / 43.3 | 38.5 / 31.7 / 30.1 | 4.8 / 3.3 / 2.9 | 8.1 / 5.8 / 4.6 | 8.2 / 5.8 / 4.1 | 22.1 / 18.2 / 16.3 |
| InternVL3-78B | 32.8 / 21.5 / 16.4 | 3.9 / 3.2 / 2.6 | 16.7 / 14.3 / 14.1 | 31.0 / 25.9 / 24.6 | 22.4 / 18.1 / 18.0 | 2.9 / 2.4 / 1.9 | 8.6 / 5.2 / 3.0 | 4.6 / 2.7 / 1.0 | 14.6 / 11.0 / 9.6 |
| InternVL3.5-8B | 24.3 / 21.6 / 19.8 | 36.6 / 35.8 / 35.6 | 33.8 / 33.3 / 33.3 | 43.9 / 43.0 / 42.1 | 40.4 / 38.3 / 37.8 | 7.6 / 6.7 / 6.7 | 44.4 / 44.1 / 43.9 | 21.5 / 20.9 / 20.5 | 33.1 / 32.0 / 31.5 |
| DeepSeek-VL2 | 10.7 / 5.5 / 1.7 | 1.0 / 1.0 / 0.3 | 6.6 / 3.8 / 1.5 | 12.9 / 9.9 / 7.6 | 7.1 / 5.3 / 5.1 | 1.9 / 1.0 / 0.0 | 2.5 / 1.1 / 0.0 | 2.6 / 1.0 / 0.0 | 5.3 / 3.3 / 1.9 |
| Qwen3-VL-8B-Instruct | 39.5 / 38.1 / 39.0 | 48.5 / 47.6 / 48.2 | 40.4 / 39.8 / 40.4 | 55.0 / 53.1 / 54.4 | 50.0 / 49.5 / 50.0 | 1.9 / 1.7 / 1.9 | 49.5 / 48.7 / 48.5 | 30.8 / 29.0 / 29.7 | 41.9 / 40.8 / 41.4 |
| Qwen3-VL-8B-Thinking | 63.3 / 63.3 / 63.3 | 58.6 / 58.6 / 58.6 | 59.1 / 58.8 / 59.1 | 67.8 / 67.7 / 67.8 | 62.8 / 62.8 / 62.8 | 9.5 / 9.0 / 8.6 | 71.2 / 71.1 / 71.2 | 52.8 / 52.6 / 52.8 | 58.2 / 58.1 / 58.1 |
| Qwen3-VL-30B-A3B-Instruct | 37.3 / 34.2 / 35.0 | 56.3 / 54.9 / 55.0 | 54.0 / 52.8 / 54.0 | 63.2 / 61.7 / 62.0 | 60.3 / 59.6 / 60.3 | 6.7 / 6.7 / 6.7 | 63.6 / 62.6 / 63.1 | 42.1 / 41.5 / 41.5 | 50.6 / 49.4 / 49.8 |
| Qwen3-VL-30B-A3B-Thinking | 68.4 / 67.8 / 67.8 | / 64.4 / 64.4 | / 67.2 / 67.2 | 75.4 / 75.1 / 74.9 | / 75.5 / 75.6 | / 11.4 / 11.4 | / 78.2 / 78.3 | / 62.4 / 62.6 | / 65.4 / 65.4 |
| Qwen3-VL-235B-A22B-Instruct | / 69.9 / 71.8 | 63.4 / 62.4 / 62.5 | 65.2 / 64.7 / 65.2 | / 75.1 / 75.4 | 71.8 / 70.9 / 71.2 | 6.7 / 6.7 / 6.7 | 69.7 / 69.2 / 69.2 | 57.4 / 56.9 / 57.4 | 63.1 / 62.2 / 62.6 |
4 Experiment
DynaSolidGeo supports the random generation of two visual versions: a randomized-view image and a 360-degree rotation video, for each question instance. Since existing geometry problem-solving tasks focus exclusively on the text-image modality, here we also evaluate models under the same text-image setting.
4.1 Experimental Setup
Evaluation Models. We evaluate a range of the latest, popular, and state-of-the-art (SOTA) closed-source and open-source MLLMs. The closed-source models include GPT-5-Nano [27], GPT-5 [27], Gemini-2.5-Flash [11], Gemini-2.5-Pro [11], and Claude-Sonnet-4.5 [5]. The open-source models include LLaVA-OneVision-1.5 family (4B, 8B) [4], GLM-4.5V [17], GLM-4.1V-9B-Thinking [17], Llama-3.2-90B-Vision-Instruct [1], Llama-4-Maverick-17B-Instruct [2], InternVL3-78B [46], InternVL3.5-8B [37], DeepSeek-VL2 [36], and the Qwen3-VL family [30].
Implementation Details. We sample batches of question instances by setting the random seed to 0, 1, and 2, respectively, resulting in a total of 1,509 text-image question instances. For answer evaluation, we allow a 1% relative error tolerance. For process evaluation, we employ Qwen3-14B [39] as the judge model. For the evaluated models, we deploy small-scale models, including the LLaVA-OneVision-1.5 family, GLM-4.1V-9B-Thinking, InternVL3.5-8B, and the Qwen3-VL family (4B, 30B), on NVIDIA A800 GPUs for evaluation. DeepSeek-VL2 is evaluated via the SiliconFlow API111https://www.siliconflow.com/, while all remaining models are accessed through the OpenRouter API222https://openrouter.ai/ for evaluation.
4.2 Experimental Results
Overall Results on Evaluation Metrics. Table 3 presents the performance of the models in Section 4.1 on the Answer Accuracy (AA), Process Score (PS), and Process-Qualified Accuracy (PA) metrics. For the GPT-5 family, LLaVA-OneVision-1.5 family, and GLM-4.5V, the PS and PA metrics are not reported, as these models either do not disclose their reasoning traces by API or inherently do not produce explicit reasoning processes. Among the closed-source models, GPT-5 achieves the highest overall AA score of 70.8%, outperforming all other models. Among the open-source models, Qwen3-VL-30B-A3B-Thinking attains the highest AA score of 65.6%, surpassing most of the closed-source models. Among the models with available reasoning traces, Qwen3-VL-30B-A3B-Thinking achieves the highest PS and PA scores, both at 65.4%.
Performance Differences across Categories from the Perspective of Spatial Intelligence. As shown in Table 3, the best-performing models perform well in Area calculation (AR), Volume calculation (VC), and Dynamic or moving-point problems (DM). However, all models struggle with Counting problems (CP), and a significant performance gap can be observed between open-source and closed-source models in this category. This discrepancy can be explained through the lens of spatial intelligence theory [14]. Tasks such as AR, VC, and DM mainly rely on lower or mid-level spatial perception, spatial relation, and spatial orientation, where visual cues are explicit and reasoning can be simplified into formula- or rule-based deduction, rather than fine-grained 3D structural reasoning. This aligns well with the representational strengths of current MLLMs, which operate on visual encodings and symbolic reasoning. In contrast, Counting Problems (CP) require higher-order mental rotation and spatial visualization, requiring 3D reconstruction, occlusion reasoning, and mental manipulation of hidden or rotated objects. This explains why models perform relatively well on AR, VC, and DM tasks but fail consistently on CP tasks.
Metric Degradation under Process Evaluation. As shown in Table 3, after introducing process evaluation, all models exhibit varying degrees of decline in both PS and PA metrics compared to AA. Among them, Gemini-2.5-Pro shows the largest drop, with PA decreasing by 9.4% relative to AA, followed by Llama-3.2-90B-Vision-Instruct, whose PA drops by 6%. This suggests that these models, while capable of producing correct answers, often rely on reasoning processes that are less coherent or causally aligned with the final answers. Furthermore, the decline in PA relative to AA is generally smaller for thinking models than for instruct models. For example, GLM-4.1V-9B-Thinking shows only a 1.5% drop, and Qwen3-VL-8B-Thinking decreases by merely 0.1%, whereas both Llama-3.2-90B-Vision-Instruct and Llama-4-Maverick-17B-Instruct experience drops exceeding 5%. Even within the Qwen3-VL family, Qwen3-VL-8B-Thinking and Qwen3-VL-30B-A3B-Thinking exhibit smaller declines compared to Qwen3-VL-8B-Instruct, Qwen3-VL-30B-A3B-Instruct, and Qwen3-VL-235B-A22B-Instruct. These observations suggest that thinking models generally produce reasoning processes that are more coherent, logically sound, and causally consistent with their final answers than those of instruct models. In addition, only a few stronger models, such as those from the Gemini and Qwen3-VL family, achieve identical AA, PS, and PA on Counting problems (CP), whereas others show clear metric gaps on different task types. This indicates that the CP task requires higher-order spatial intelligence and more rigorous symbolic reasoning. Consequently, making correct answers less susceptible to hallucination or logical inconsistency.
Data Contamination and Memorization Phenomenon. To probe potential data contamination and memorization effects of VLMs on static datasets, we further evaluate their Answer Accuracy (AA) on the static source questions, as shown in Figure 3. Compared with the static source-question dataset, all models show a notable performance drop on DynaSolidGeo, with Claude-Sonnet-4.5 (-20.4%) and InternVL3.5-8B (-17.6%) declining the most. This reveals that these VLMs may suffer from varying degrees of data contamination on static datasets and tend to rely on memorization-based patterns rather than genuine reasoning processes when producing answers. In contrast, DynaSolidGeo serves as a benchmark for assessing the VLMs’ genuine ability in spatial mathematical reasoning. Furthermore, thinking models show smaller performance declines than instruct ones, suggesting that RL promotes reasoning-oriented behavior, while SFT encourages memorization of answers.
| Model | Num. of Output Tokens | Metrics |
| Correct / Incorrect / All | AA / PS / PA | |
| Closed-sourced MLLMs | ||
| GPT-5-Nano | 8846.12 / 9739.68 / 9280.17 | 51.4 / - / - |
| GPT-5 | 7958.86 / 9900.60 / 8526.33 | 70.8 / - / - |
| Gemini-2.5-Flash | 4865.65 / / 15231.90 | 49.6 / 47.8 / 47.9 |
| Gemini-2.5-Pro | / 18663.74 / | 62.0 / 55.9 / 52.6 |
| Claude-Sonnet-4.5 | 987.64 / 1047.81 / 1028.15 | 32.7 / 29.7 / 28.6 |
| Open-sourced VLMs | ||
| LLaVA-OneVision-1.5-4B-Instruct | 712.62 / 842.79 / 833.90 | 6.8 / - / - |
| LLaVA-OneVision-1.5-8B-Instruct | 262.11 / 150.33 / 156.19 | 5.2 / - / - |
| GLM-4.5V | 4679.18 / 5287.26 / 5055.40 | 38.1 / - / - |
| GLM-4.1V-9B-Thinking | 6599.09 / 7562.91 / 7309.98 | 26.2 / 24.6 / 24.7 |
| Llama-3.2-90B-Vision-Instruct | 756.23 / 821.67 / 806.88 | 22.6 / 18.5 / 16.6 |
| Llama-4-Maverick-17B-Instruct | 761.78 / 866.27 / 843.14 | 22.1 / 18.2 / 16.3 |
| InternVL3-78B | 562.24 / 599.96 / 594.46 | 14.6 / 11.0 / 9.6 |
| InternVL3.5-8B | 11336.19 / 15019.64 / 13801.59 | 33.1 / 32.0 / 31.5 |
| DeepSeek-VL2 | 404.41 / 465.09 / 461.88 | 5.3 / 3.3 / 1.9 |
| Qwen3-VL-8B-Instruct | 7514.76 / 18246.12 / 13751.61 | 41.9 / 40.8 / 41.4 |
| Qwen3-VL-8B-Thinking | / 21482.31 / | 58.2 / 58.1 / 58.1 |
| Qwen3-VL-30B-A3B-Instruct | 10236.35 / / 16077.87 | 50.6 / 49.4 / 49.8 |
| Qwen3-VL-30B-A3B-Thinking | 10954.33 / 17192.55 / 13099.88 | 65.6 / 65.4 / 65.4 |
| Qwen3-VL-235B-A22B-Instruct | 6045.59 / 6945.21 / 6370.42 | 63.1 / 62.2 / 62.6 |
Model Inference Efficiency Analysis. Table 4 summarizes, for each model, the average number of output tokens in the cases of correct and incorrect answers, as well as the overall average, together with the corresponding performance metrics. Overall, the number of output tokens is roughly positively correlated with model performance—models with very poor accuracy consistently produce shorter outputs, which aligns with the principle of test-time scaling. In addition, for almost all models (except LLaVA-OneVision-1.5-8B-Instruct), the reasoning traces for incorrect answers are noticeably longer than those for correct ones. This is likely because, when a model encounters a problem it cannot solve or faces logical inconsistencies during reasoning, it tends to repeatedly “rethink” its intermediate steps, resulting in unnecessarily prolonged reasoning chains.
4.3 Error Analysis
We conduct an error analysis on two representative models, Gemini-2.5-Pro (closed source) and Qwen3-VL-30B-A3B-Thinking (open source). Specifically, we categorize the errors in Process-Qualified Accuracy (PA) into five types: Visual Perception Errors, Logical Reasoning Errors, Calculation Errors, Knowledge Errors, and Hallucination Errors, as shown in Figure 4. Across the 1,509 sampled instances, Gemini-2.5-Pro makes a total of 715 errors, Qwen3-VL-30B-A3B-Thinking makes 525 errors. Among these error types, visual perception and logical reasoning errors dominate, with visual perception errors accounting for the largest proportion. This indicates that, although these advanced models have demonstrated strong symbolic reasoning capabilities, they still lack sufficient perceptual understanding in solid geometry tasks that require spatial intelligence. Qwen3-VL-30B-A3B-Thinking exhibits 52 more Visual Perception errors than Gemini-2.5-Pro, suggesting that it is more prone to reasoning failures triggered by inaccurate visual perception. This also explains why Qwen3-VL-30B-A3B-Thinking performs significantly worse than Gemini-2.5-Pro on Counting Problems (CP), which demand higher levels of spatial intelligence. In addition, Qwen3-VL-30B-A3B-Thinking makes fewer errors in other categories (including Logical Reasoning Errors), demonstrating its stronger symbolic reasoning capability.
5 Conclusion
In this work, we introduced DynaSolidGeo, the first dynamic benchmark for evaluating the genuine spatial mathematical reasoning capabilities of VLMs in solid geometry. Through a semi-automatic, expert-guided pipeline, DynaSolidGeo enables unbounded generation of diverse multimodal instances, effectively mitigating contamination and memorization issues found in static datasets. By integrating both answer- and process-level evaluation, we provide a more faithful assessment of reasoning validity and spatial perception. Comprehensive experiments uncover persistent limitations in high-level spatial intelligence and reveal substantial performance degradation under dynamic evaluation. We expect DynaSolidGeo to provide a reliable foundation for advancing process-grounded, contamination-resistant multimodal reasoning benchmarks and inspire future research toward robust spatial reasoning in VLMs.
Future Work
We will update our Appendix in the next version.

### v2

@@
 | MathVerse [58] | EN | 119 (15.1%) | - | K. | Static | Text&Image | A.E. |
+| GeomRel [42] | EN | 0 (0.0%) | - | K. | Static | Text&Image | A.E.&P.E. |
 | GeoSense [50] | EN&CN | 350 (20.0%) | 2 | K. | Static | Text&Image | A.E.&P.E. |
@@
 1 Introduction
-Geometry problem solving has long played a central role in mathematical reasoning, requiring integrating visual understanding and symbolic reasoning across complex graphic and textual contexts [38].
+Geometry problem solving has long played a central role in mathematical reasoning, requiring integrating visual understanding and symbolic reasoning across complex graphic and textual contexts [48, 50, 42, 63, 35, 32, 44, 30, 43, 10, 21].
 According to structural properties, geometry can be categorized into plane geometry and solid geometry.
-Compared to plane geometry, solid geometry imposes substantially higher spatial mathematical reasoning ability, as reasoning in three dimensions entails spatial intelligence, including spatial perception, spatial relation, spatial orientation, spatial rotation, and spatial visualization that goes beyond two-dimensional recognition [14, 29, 16, 22].
-Such tasks remain difficult even for human learners [7], and represent a formidable open challenge for current AI systems.
+Compared to plane geometry, solid geometry imposes substantially higher spatial mathematical reasoning ability, as reasoning in three dimensions entails spatial intelligence, including spatial perception, spatial relation, spatial orientation, spatial rotation, and spatial visualization that goes beyond two-dimensional recognition [51, 53, 16, 36, 19, 26].
+Such tasks remain difficult even for human learners [7], and represent a formidable open challenge for current AI systems [59].
 In parallel, recent years have witnessed remarkable progress in multimodal large language models (MLLMs).
-Building on the successes of foundation models, vision-language models (VLMs) [23, 6, 11, 27, 5, 4, 17] have rapidly advanced the state of the art in a wide spectrum of multimodal understanding tasks.
-Among these tasks, multimodal mathematical reasoning has emerged as a challenging yet vibrant frontier, with benchmarks such as GeoQA [8], MathVista [25], and GeoSense [38] exposing both the promise and the limitations of current VLMs.
+Building on the successes of foundation models, vision-language models (VLMs) [27, 6, 12, 33, 5, 4, 20, 14, 17, 60] have rapidly advanced the state of the art in a wide spectrum of multimodal understanding tasks.
+Among these tasks, multimodal mathematical reasoning has emerged as a challenging yet vibrant frontier, with benchmarks such as GeoQA [8], MathVista [29], GeomRel [42], and GeoSense [50] exposing both the promise and the limitations of current VLMs.
 These carefully designed benchmarks have played a pivotal role in advancing the field, providing standardized evaluation and catalyzing iterative improvements in model design and training paradigm.
@@
 Second, nearly all existing multimodal mathematical reasoning benchmarks are static, relying on fixed and finite test sets that are susceptible to data contamination and memorization.
-Recent analyses demonstrate that large models can memorize and regurgitate benchmark data [26, 12, 21, 28, 10], and some studies show that decontaminated re-releases often lead to substantial drops in performance [44], confirming that static evaluation may significantly overestimate true reasoning and generalization [31].
-Similar concerns have motivated dynamic evaluations in coding [18, 45] and general-purpose QA [34, 35], indicating a community-wide shift towards dynamic, contamination-resistant evaluation protocols [9].
+Recent analyses demonstrate that large models can memorize and regurgitate benchmark data [31, 13, 25, 34, 11, 37], and some studies show that decontaminated re-releases often lead to substantial drops in performance [62], confirming that static evaluation may significantly overestimate true reasoning and generalization [39].
+Similar concerns have motivated dynamic evaluations in coding [22, 64] and general-purpose QA [45, 46], indicating a community-wide shift towards dynamic, contamination-resistant evaluation protocols [9, 61].
 Moreover, most existing multimodal mathematical reasoning benchmarks [29, 41, 40, 56, 55, 57, 66] evaluate models solely based on answer accuracy, which allows models suffering from data contamination or over-reliance on memorization to appear strong, while failing to reveal their genuine reasoning ability.
@@
 Unlike existing static resources, DynaSolidGeo consists of 503 seed questions of solid geometry problem solving, each represented by a Python program paired with a corresponding MATLAB program.
-With the correctness of the question guaranteed, each seed question is parameterized: textual variables in the question statement (e.g., endpoint labels, side lengths, areas, volumes, ratio) as well as rendering parameters of the solid geometry (e.g., camera viewpoints) can all be randomized.
-By supplying different random seeds, DynaSolidGeo can in principle generate an unbounded number of question-answer instances, where each instance can optionally include two visual versions: a randomized-view image and a 360-degree rotation video.
+With the correctness of the question guaranteed, each seed question is parameterized: textual variables in the question statement (e.g., endpoint labels, side lengths, areas, volumes, ratios) as well as rendering parameters of the solid geometry (e.g., camera viewpoints) can all be randomized.
+By supplying different random seeds, DynaSolidGeo can, in principle, generate an unbounded number of question-answer instances, where each instance can optionally include two visual versions: a randomized-view image and a 360-degree rotation video.
 The seed questions of DynaSolidGeo are drawn from diverse and authoritative sources, including China’s Gaokao examinations, international mathematics competitions, and widely used training materials for competition preparation.
@@
 Early efforts include TQA [23] and Geometry3K [28] introduced multimodal reasoning tasks involving diagram-based science and geometry word problems with accompanying 2D visuals.
-More recent benchmarks, such as GeoQA [8], PGPS9K [42], MMMU-MATH [40], GeoEval [41], MATH-Vision [32], OlympiadBench [15], MathVerse [43], GeoSense [38], and GeoLaux [13] have broadened coverage to thousands of multimodal math problems.
-However, the vast majority of these resources concentrate on plane geometry and 2D diagrammatic reasoning, leaving solid geometry largely underexplored.
+More recent benchmarks, such as GeoQA [8], PGPS9K [57], MMMU-MATH [55], GeoEval [56], MATH-Vision [40], OlympiadBench [18], MathVerse [58], GeomRel [42], GeoSense [50], and GeoLaux [15] have broadened coverage to thousands of multimodal math problems.
+However, most of these resources focus on plane geometry and 2D diagrammatic reasoning, leaving solid geometry largely underexplored.
 A few datasets have attempted to move toward 3D: SolidGeo [41] explicitly targets solid geometry but remains static datasets vulnerable to contamination and memorization; DynaMath [66] introduces dynamic instance generation, but solid geometry is barely represented, with only 15 problems (3%) in the dataset.
@@
 In parallel, the open-source community has introduced competitive alternatives such as LLaVA-OneVision-1.5 family [4], GLM-4.1V-9B-Thinking [20], Llama-4-Maverick-17B-Instruct [2], InternVL3.5-8B [49], DeepSeek-VL2 [47], and the Qwen3-VL family [38], which push the frontier of visual capabilities.
-Yet their spatial mathematical reasoning ability remains underexplored, motivating our evaluation on DynaSolidGeo.
+Yet their spatial mathematical reasoning ability remains underexplored, motivating our evaluation of DynaSolidGeo.
 3 DynaSolidGeo
@@
 1.
-Expert-Guided Parametrization and Visualization: Here we follows a human-in-the-loop strategy, where human experts collaborate with large models to create a JSON annotation and a MATLAB program for each source question:
+Expert-Guided Parametrization and Visualization: Here we follow a human-in-the-loop strategy, where human experts collaborate with large models to create a JSON annotation and a MATLAB program for each source question:
 JSON-Based Question Parametrization: For each collected source question, mathematics experts parameterize the question statement by converting fixed values into variable parameters (e.g., endpoint labels, side lengths, areas, volumes, ratios) using f-string syntax, while ensuring correctness and availability.
-Corresponding answer is also expressed in terms of these variable parameters.
+The corresponding answer is also expressed in terms of these variable parameters.
 An example of parameterized variables is highlighted in red in Fig.
@@
 Table 2 summarizes the detailed statistics of the DynaSolidGeo dataset.
-In total, the benchmark contains 503 curated seed questions, all newly constructed for this work.
+In total, the benchmark contains 503 curated seed questions, all newly constructed for this work, and provides both Chinese and English versions for each question statement.
 The questions span a diverse set of solid geometry problem categories, including Positional relationship determination (PD, 11.7%), Angle calculation (AN, 20.5%), Length and distance calculation (LC, 13.1%), Area calculation (AR, 11.3%), Volume calculation (VC, 10.3%), Counting problems (CP, 7.0%), Dynamic or moving-point problems (DM, 13.1%), and Folding and unfolding problems (FP, 12.9%).
@@
 | - Newly curated questions | 503 (100.0%) |
+| - English ver. / Chinese ver. | 503 (100.0%) / 503 (100.0%) |
 | Categories | |
@@
 3.4.2 Process Score
-To more accurately assess the quality of the model’s reasoning process, we annotate not only parameterized question statements and answers, but also parameterized reference reasoning processes.
-Each reasoning process is evaluated using LLM as a judge according to the following criteria:
+To more accurately assess the quality of the reasoning process, we also annotate parameterized reference reasoning chains.
+Based on these expert-annotated reasoning chains, each reasoning process is evaluated using LLM as a judge according to the following criteria:
 Logical Alignment: The reasoning presents a coherent derivation whose steps consistently lead to the stated result, with matching variables/units and without any unjustified conclusion jumps.
@@
 Since existing geometry problem-solving tasks focus exclusively on the text-image modality, here we also evaluate models under the same text-image setting.
+The performance of some VLMs in the text-video modality can be found in Appendix D.2.
 4.1 Experimental Setup
@@
 DeepSeek-VL2 is evaluated via the SiliconFlow API111https://www.siliconflow.com/, while all remaining models are accessed through the OpenRouter API222https://openrouter.ai/ for evaluation.
+We set the temperature to 0.0 for all models to reduce randomness, while keeping all other hyperparameters at their default values.
+More details of the experiment setup can be found in Appendix C.
 4.2 Experimental Results
@@
 However, all models struggle with Counting problems (CP), and a significant performance gap can be observed between open-source and closed-source models in this category.
-This discrepancy can be explained through the lens of spatial intelligence theory [14].
+This discrepancy can be explained through the lens of spatial intelligence theory [16] (for more background on spatial intelligence, see Appendix B).
 Tasks such as AR, VC, and DM mainly rely on lower or mid-level spatial perception, spatial relation, and spatial orientation, where visual cues are explicit and reasoning can be simplified into formula- or rule-based deduction, rather than fine-grained 3D structural reasoning.
@@
 In addition, only a few stronger models, such as those from the Gemini and Qwen3-VL family, achieve identical AA, PS, and PA on Counting problems (CP), whereas others show clear metric gaps on different task types.
-This indicates that the CP task requires higher-order spatial intelligence and more rigorous symbolic reasoning.
-Consequently, making correct answers less susceptible to hallucination or logical inconsistency.
+This indicates that the CP task requires higher-order spatial intelligence and more rigorous symbolic reasoning, consequently making correct answers less susceptible to hallucination or logical inconsistency.
 Data Contamination and Memorization Phenomenon.
@@
 In addition, Qwen3-VL-30B-A3B-Thinking makes fewer errors in other categories (including Logical Reasoning Errors), demonstrating its stronger symbolic reasoning capability.
+More details on the error analysis can be found in Appendix D.3.
 5 Conclusion
@@
 Through a semi-automatic, expert-guided pipeline, DynaSolidGeo enables unbounded generation of diverse multimodal instances, effectively mitigating contamination and memorization issues found in static datasets.
-By integrating both answer- and process-level evaluation, we provide a more faithful assessment of reasoning validity and spatial perception.
+By integrating both answer- and process-level evaluation, we provide a more faithful assessment of logical validity and causal coherence.
 Comprehensive experiments uncover persistent limitations in high-level spatial intelligence and reveal substantial performance degradation under dynamic evaluation.
-We expect DynaSolidGeo to provide a reliable foundation for advancing process-grounded, contamination-resistant multimodal reasoning benchmarks and inspire future research toward robust spatial reasoning in VLMs.
-Future Work
-We will update our Appendix in the next version.
+We expect DynaSolidGeo to provide a reliable foundation for advancing process-grounded multimodal reasoning benchmarks that mitigate data contamination and inspire future research toward robust spatial reasoning in VLMs.
+Supplementary Material
+Appendix A Limitations and Future Work
+First, DynaSolidGeo provides both Chinese and English versions of question statements.
+All evaluations in this paper are conducted on the English version, while model performance on the Chinese version remains underexplored.
+Future work can investigate how VLMs perform on the Chinese version of DynaSolidGeo.
+Second, although our process evaluation combines expert-annotated reasoning chains with an LLM-as-a-judge approach, where the former mitigates bias from different judge models and the latter alleviates the limitations of single-path expert annotations, the gap between this hybrid evaluation method and full human evaluation has not been thoroughly examined.
+Exploring more robust and practical process evaluation methods thus represents an essential step toward reliable and comprehensive large-model assessment.
+Appendix B Background of Spatial Intelligence
+Spatial intelligence [51, 53, 16, 36, 19, 26, 54] refers to the cognitive ability to perceive, understand, and mentally manipulate spatial relationships among objects.
+In cognitive science, it is typically decomposed into five core dimensions [51, 53, 16, 36, 19, 26]:
+Spatial Perception: The ability to accurately perceive horizontal and vertical orientations while resisting misleading visual context.
+It supports basic posture stabilization, balance maintenance, and spatial alignment.
+Spatial Relation: The ability to rapidly infer positional relationships or assembly structures between multiple components, often realized through mental rotation of simple shapes for quick matching.
+Spatial Orientation: The ability to maintain correct directional judgments after viewpoint transformations, such as imagining oneself moving to another location and identifying the relative direction of objects.
+Mental Rotation: The ability to mentally rotate 2D or 3D objects continuously and match them with target configurations.
+Spatial Visualization: The capability to perform multi-step and complex mental transformations such as folding, cutting, or assembling objects, enabling transitions between 2D and 3D representations.
+These five dimensions reflect an increasing hierarchy of spatial cognition, from basic perception to high-order 3D manipulation, and form the theoretical foundation for evaluating spatial mathematical reasoning in solid geometry.
+In DynaSolidGeo, different task categories depend on spatial intelligence to varying degrees (e.g., AR/VC/DM largely rely on lower- to mid-level abilities such as perception and relational reasoning, whereas CP strongly requires high-level mental rotation and spatial visualization), enabling us to interpret performance differences across VLMs better.
+Appendix C More Details of Experiment Setup
+In this section, we provide more details of the experiment setup.
+C.1 Prompt Template for Response Generation
+[…]
+C.2 Prompt Template for Judge Model
+For process evaluation, we adopt an LLM-as-a-judge approach, where a LLM (Qwen3-14B) assesses the reasoning process of each evaluated model against our expert-annotated reference reasoning chain.
+The judge model is prompted to output one of the discrete scores 0, 0.25, 0.5, 0.75, 1 as the Process Score (PS), where 0 denotes “Not Supportive/Incorrect”, 0.25 denotes “Marginally Related”, 0.5 denotes “Partial Support”, 0.75 denotes “Near Causality”, and 1 denotes “Full Causality”.
+The prompt template is shown below.
+C.3 Model Hyperparameters
+We set the temperature to 0.0 for all models to reduce randomness, while keeping all other hyperparameters at their default values. Table 5 shows the model hyperparameters.
+| Model | Hyperparameters |
+| GPT-5-Nano | model = openai/gpt-5-nano, temperature = 0.0, max tokens = 128K |
+| GPT-5 | model = openai/gpt-5, temperature = 0.0, max tokens = 128K |
+| Gemini-2.5-Flash | model = google/gemini-2.5-flash, temperature = 0.0, max tokens = 65.5K |
+| Gemini-2.5-Pro | model = google/gemini-2.5-pro, temperature = 0.0, max tokens = 65.5K |
+| Claude-Sonnet-4.5 | model = anthropic/claude-sonnet-4.5, temperature = 0.0, max tokens = 64K |
+| LLaVA-OneVision-1.5-4B-Instruct | model = lmms-lab/LLaVA-OneVision-1.5-4B-Instruct, temperature = 0.0, max tokens = 1024 |
+| LLaVA-OneVision-1.5-8B-Instruct | model = lmms-lab/LLaVA-OneVision-1.5-8B-Instruct, temperature = 0.0, max tokens = 1024 |
+| GLM-4.5V | model = z-ai/glm-4.5v, temperature = 0.0, max tokens = 16.4K |
+| GLM-4.1V-9B-Thinking | model = thudm/glm-4.1v-9b-thinking, temperature = 0.0, max tokens = 8K |
+| Llama-3.2-90B-Vision-Instruct | model = meta-llama/llama-3.2-90b-vision-instruct, temperature = 0.0, max tokens = 32.8K |
+| Llama-4-Maverick-17B-Instruct | model = meta-llama/llama-4-maverick, temperature = 0.0, max tokens = 16.4K |
+| InternVL3-78B | model = opengvlab/internvl3-78b, temperature = 0.0, max tokens = 32.8K |
+| InternVL3.5-8B | model = internlm/CapRL-InternVL3.5-8B, temperature = 0.0, max tokens = 32K |
+| DeepSeek-VL2 | model = deepseek-ai/deepseek-vl2, temperature = 0.0, max tokens = 128K |
+| Qwen3-VL-8B-Instruct | model = Qwen/Qwen3-VL-8B-Instruct, temperature = 0.0, max tokens = 32K |
+| Qwen3-VL-8B-Thinking | model = Qwen/Qwen3-VL-8B-Thinking, temperature = 0.0, max tokens = 32K |
+| Qwen3-VL-30B-A3B-Instruct | model = Qwen/Qwen3-VL-30B-A3B-Instruct, temperature = 0.0, max tokens = 32K |
+| Qwen3-VL-30B-A3B-Thinking | model = Qwen/Qwen3-VL-30B-A3B-Thinking, temperature = 0.0, max tokens = 32K |
+| Qwen3-VL-235B-A22B-Instruct | model = Qwen/Qwen3-VL-235B-A22B-Instruct, temperature = 0.0, max tokens = 32K |
+| Model | PD | AN | LC | AR | VC | CP | DM | FP | ALL |
+| AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | |
+| Closed-sourced MLLMs | |||||||||
+| GPT-5-Nano | 40.7 / - / - | 53.4 / - / - | 54.5 / - / - | 77.2 / - / - | 69.2 / - / - | 5.7 / - / - | 56.1 / - / - | 36.9 / - / - | 51.3 / - / - |
+| GPT-5 | 76.3 / - / - | 65.0 / - / - | 78.8 / - / - | 80.7 / - / - | 84.6 / - / - | 17.1 / - / - | 75.8 / - / - | 63.1/ - / - | 69.8 / - / - |
+| Gemini-2.5-Flash | 44.1 / 44.1 / 44.1 | 48.5/ 47.6/ 47.6 | 60.6 / 58.7/ 59.1 | 61.0 / 57.9 / 57.9 | 55.6 / 51.1 / 51.1 | 17.1 / 17.1 / 17.1 | 72.7 / 71.1 / 72.7 | 30.1 / 30.0 / 30.1 | 50.5 / 49.1 / 49.3 |
+| Gemini-2.5-Pro | 72.1 / 67.4 / 66.1 | 52.4 / 45/ 41.8 | 72.7 / 68.9 / 68.2 | 65.0 / 61.4 / 59.7 | 75.0 / 66.4 / 59.6 | 45.7 / 45.7 / 45.7 | 62.1 / 56.8 / 53.0 | 56.9 / 51.9 / 50.8 | 62.6 / 57.3 / 54.9 |
+| Claude-Sonnet-4.5 | 49.2 / 40.7 / 37.3 | 25.2 / 23.5 / 22.3 | 40.9 / 37.5 / 34.9 | 54.4 / 51.2 / 52.6 | 57.7 / 53.9 / 50.0 | 8.6 / 7.9 / 8.6 | 21.2 / 20.5 / 19.7 | 16.9 / 16.2 / 16.9 | 34.0 / 31.3 / 30.0 |
+| Open-sourced VLMs | |||||||||
+| LLaVA-OneVision-1.5-4B-Instruct | 13.6 / - / - | 2.9 / - / - | 6.1 / - / - | 15.8 / - / - | 11.5 / - / - | 2.9 / - / - | 1.5 / - / - | 0.0 / - / - | 6.4 / - / - |
+| LLaVA-OneVision-1.5-8B-Instruct | 22.0 / - / - | 1.0 / - / - | 7.6 / - / - | 3.5 / - / - | 3.9 / - / - | 0.0 / - / - | 6.1 / - / - | 4.6 / - / - | 6.0 / - / - |
+| GLM-4.5V | 54.2 / - / - | 32/ - / - | 45.5 / - / - | 56.1 / - / - | 44.2 / - / - | 5.7 / - / - | 45.5/ - / - | 13.9 / - / - | 38.0 / - / - |
+| GLM-4.1V-9B-Thinking | 23.7 / 20.8 / 18.6 | 22.3 / 20.9 / 21.4 | 28.8 / 27.7 / 28.8 | 45.6 / 44.3 / 43.9 | 34.6 / 32.2 / 32.7 | 2.9 / 2.9 / 2.9 | 21.2 / 20.5 / 19.7 | 4.6 / 3.9 / 3.1 | 23.5 / 22.1 / 21.9 |
+| Llama-3.2-90B-Vision-Instruct | 33.9 / 27.1 / 23.7 | 9.7 / 8.5 / 7.8 | 28.8 / 26.5 / 25.8 | 54.4 / 48.7 / 43.4 | 40.4 / 36.5 / 34.6 | 2.9 / 1.4 / 0.0 | 9.1 / 6.4 / 6.1 | 7.7 / 3.9 / 1.5 | 22.5 / 19.1 / 17.7 |
+| Llama-4-Maverick-17B-Instruct | 37.3 / 30.9 / 30.5 | 15.5 / 12.1 / 9.7 | 25.8 / 22.7 / 22.7 | 50.9 / 46.9 / 43.9 | 40.4 / 31.7 / 30.8 | 5.7 / 2.9 / 2.9 | 12.1 / 9.5 / 7.6 | 10.8 / 8.5 / 7.7 | 24.3 / 20.2 / 18.9 |
+| InternVL3-78B | 3.0 / 23.7 / 18.6 | 3.9 / 3.4 / 2.9 | 16.7 / 13.6 / 13.6 | 29.8 / 23.3 / 21.1 | 25.0 / 19.7 / 21.2 | 5.7 / 5.7 / 5.7 | 6.1 / 3.4 / 1.5 | 3.8 / 2.3 / 1.5 | 14.5 / 11.1 / 9.9 |
+| InternVL3.5-8B | 23.7 / 20.3 / 18.6 | 36.9 / 35.4 / 35.0 | 33.3 / 33.0 / 33.3 | 45.6 / 44.7 / 43.9 | 38.5 / 36.5 / 36.5 | 8.6 / 7.1 / 8.6 | 45.5 / 45.5 / 45.5 | 15.4 / 15.4 / 15.4 | 32.4 / 31.3 / 31.0 |
+| DeepSeek-VL2 | 13.6 / 7.2 / 5.1 | 1.9 / 0.5 / 0.0 | 7.6 / 4.9 / 3.0 | 14.0 / 11.4 / 8.8 | 5.8 / 5.3 / 5.8 | 0.0 / 0.0 / 0.0 | 0.0 / 0.0 / 0.0 | 0.0 / 0.0 / 0.0 | 5.2 / 3.4 / 2.6 |
+| Qwen3-VL-8B-Instruct | 50.9 / 49.2 / 50.9 | 52.4 / 51.9 / 52.4 | 42.4 / 42.4 / 42.4 | 61.4 / 59.7 / 61.4 | 59.6 / 59.1 / 59.6 | 2.9 / 2.1 / 2.9 | 5.0 / 50.0 / 50.0 | 32.3 / 31.2 / 30.8 | 46.3 / 45.6 / 46.1 |
+| Qwen3-VL-8B-Thinking | 35.6 / 35.6 / 35.6 | 55.3 / 53.0 / 53.4 | 56.1 / 54.1 / 56.1 | 57.9 / 57.5 / 57.9 | 59.2 / 59.6 / 59.6 | 5.7 / 5.7 / 5.7 | 65.2 / 64.0/ 65.2 | 40.0 / 38.9 /38.5 | 49.7 / 48.8 / 49.1 |
+| Qwen3-VL-30B-A3B-Instruct | 37.3 / 34.2 / 35.0 | 56.3 / 54.9 / 55.0 | 54.0 / 52.8 / 54.0 | 63.2 / 61.7 / 62.0 | 60.3 / 59.6 / 60.3 | 6.7 / 6.7 / 6.7 | 63.6 / 62.6 / 63.1 | 42.1 / 41.5 / 41.5 | 50.6 / 49.4 / 49.8 |
+| Qwen3-VL-30B-A3B-Thinking | 67.8 / 67.8 / 67.8 | 63.1 / 63.1 / 63.1 | 69.7 / 69.7 / 69.7 | 79.0 / 79.0 / 79.0 | 71.2 / 69.2 / 69.2 | 11.4 / 11.4 / 11.4 | 75.8 / 75.8 / 75.8 | 60.0 / 59.6 / 60.0 | 64.8 / 64.6 / 64.6 |
+| Qwen3-VL-235B-A22B-Instruct | 67.8 / 66.1 / 67.8 | 63.1 / 62.4 / 62.1 | 72.7 / 72.3 / 72.7 | 73.7 / 72.4 / 71.9 | 69.2 / 68.8 / 69.2 | 5.7 / 5.7 / 5.7 | 74.2 / 74.2 / 74.2 | 56.9 / 56.2 / 56.9 | 63.4 / 62.7 / 63.0 |
+| Model | PD | AN | LC | AR | VC | CP | DM | FP | ALL |
+| AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | |
+| Closed-sourced MLLMs | |||||||||
+| GPT-5-Na no | 40.7 / - / - | 52.4 / - / - | 48.5 / - / - | 70.2 / - / - | 73.1 / - / - | 5.7 / - / - | 47.0 / - / - | 46.2 / - / - | 49.9 / - / - |
+| GPT-5 | 71.2 / - / - | 68.9 / - / - | 80.3 / - / - | 87.7 / - / - | 84.6 / - / - | 17.1 / - / - | 78.8 / - / - | 67.7 / - / - | 72.0 / - / - |
+| Gemini-2.5-Flash | 44.1 / 42.4 / 42.4 | 47.6 / 45.2 / 44.7 | 59.1 / 54.2 / 54.6 | 63.2 / 59.2 / 59.7 | 53.9 / 53.9 / 53.9 | 22.9 / 22.9 / 22.9 | 63.6 / 60.1 / 60.6 | 44.6 / 41.9 / 41.5 | 51.1 / 48.6 / 48.5 |
+| Gemini-2.5-Pro | 62.7 / 49.2 / 39.0 | 50.4 / 40.3 / 34.0 | 68.2 / 59.9 / 54.6 | 73.7 / 69.3 / 66.7 | 76.9 / 69.7 / 65.4 | 31.4 / 31.4 / 31.4 | 63.6 / 61.7 / 60.6 | 52.3 / 45.0 / 40.0 | 60.2 / 53.0 / 48.3 |
+| Claude-Sonnet-4.5 | 37.3 / 30.9 / 28.8 | 25.2 / 23.5 / 22.3 | 36.4 / 33.0 / 31.8 | 45.6 / 45.2 / 45.6 | 50.0 / 45.2 / 42.3 | 8.6 / 7.2 / 5.7 | 24.2 / 22.7 / 22.7 | 15.4 / 13.9 / 12.3 | 30.4 / 27.8 / 26.64 |
+| Open-sourced VLMs | |||||||||
+| LLaVA-OneVision-1.5-4B-Instruct | 15.3 / - / - | 5.8 / - / - | 6.1 / - / - | 19.3 / - / - | 13.5 / - / - | 2.9 / - / - | 3.0 / - / - | 0.0 / - / - | 8.0 / - / - |
+| LLaVA-OneVision-1.5-8B-Instruct | 15.3 / - / - | 2.9 / - / - | 7.6 / - / - | 3.5 / - / - | 1.9 / - / - | 2.9 / - / - | 6.1 / - / - | 1.5 / - / - | 5.2 / - / - |
+| GLM-4.5V | 40.7 / - / - | 34.0 / - / - | 47.0 / - / - | 56.1 / - / - | 55.8 / - / - | 11.4 / - / - | 43.9 / - / - | 18.5 / - / - | 39.0 / - / - |
+| GLM-4.1V-9B-Thinking | 30.5 / 28.4 / 28.8 | 27.2 / 26.2 / 26.2 | 37.9 / 36.0 / 37.9 | 45.6 / 43.9 / 43.9 | 50.0 / 46.2 / 48.1 | 5.7 / 5.7 / 5.7 | 28.8 / 28.0 / 27.3 | 4.6 / 3.9 / 4.6 | 29.2 / 27.7 / 28.3 |
+| Llama-3.2-90B-Vision-Instruct | 33.9 / 25.0 / 22.0 | 17.5 / 12.4 / 9.7 | 20.8 / 19.7 / 47.4 | 47.4 / 43.4 / 40.4 | 40.4 / 31.2 / 26.9 | 2.9 / 2.1 / 2.9 | 7.6 / 6.4 / 6.1 | 6.2 / 3.9 / 1.5 | 22.3 / 17.8 / 15.7 |
+| Llama-4-Maverick-17B-Instruct | 39.0 / 27.5 / 18.6 | 14.6 / 10.1 / 6.8 | 25.8 / 22.7 / 21.2 | 45.6 / 44.3 / 43.9 | 36.5 / 30.3 / 28.9 | 8.6 / 7.1 / 5.7 | 6.1 / 3.8 / 3.0 | 3.1 / 1.5 / 0.0 | 21.7 / 17.6 / 15.1 |
+| InternVL3-78B | 30.5 / 20.8 / 15.2 | 3.9 / 2.7 / 1.9 | 16.7 / 14.0 / 12.1 | 29.8 / 26.8 / 28.1 | 23.1 / 17.3 / 15.4 | 0.0 / 0.0 / 0.0 | 10.6 / 6.8 / 4.6 | 3.1 / 1.9 / 1.5 | 14.1 / 10.8 / 9.3 |
+| InternVL3.5-8B | 22.0 / 19.9 / 18.6 | 36.9 / 36.9 / 36.9 | 34.9 / 33.7 / 33.3 | 45.6 / 44.7 / 43.9 | 44.2 / 41.4 / 40.4 | 8.6 / 7.1 / 5.7 | 51.5 / 51.1 / 51.5 | 26.2 / 25.4 / 24.6 | 35.2 / 34.1 / 33.6 |
+| DeepSeek-VL2 | 6.8 / 3.4 / 0.0 | 0.0 / 0.0 / 0.0 | 4.6 / 2.3 / 0.0 | 14.0 / 10.1 / 7.0 | 11.5 / 7.2 / 5.8 | 2.9 / 1.4 / 0.0 | 4.5 / 1.9 / 0.0 | 1.5 / 0.1 / 0.0 | 5.2 / 3.0 / 1.4 |
+| Qwen3-VL-8B-Instruct | 33.9 / 33.9 / 33.9 | 45.6 / 44.4 / 45.6 | 40.9 / 40.1 / 40.9 | 52.6 / 51.3 / 52.6 | 51.9 / 51.0 / 51.9 | 0.0 / 0.0 / 0.0 | 4.9 / 42.8 / 42.4 | 30.8 / 28.1 / 29.2 | 39.8 / 38.7 / 39.4 |
+| Qwen3-VL-8B-Thinking | 62.7 / 62.7 / 62.7 | 59.2 / 59.2 / 59.2 | 60.6 / 60.2 / 60.6 | 70.2 / 70.2 / 70.2 | 67.3 / 67.3 / 67.3 | 5.7 / 5.7 / 5.7 | 66.7 / 66.7 / 66.7 | 63.1 / 62.7 / 63.1 | 59.6 / 59.5 / 59.6 |
+| Qwen3-VL-30B-A3B-Instruct | 40.7 / 34.3 / 35.6 | 54.4 / 52.9 / 52.4 | 54.6 / 53.0 / 54.6 | 68.4 / 67.5 / 68.4 | 63.5 / 62.0 / 63.5 | 8.6 / 8.6 / 8.6 | 63.6 / 62.5 / 62.1 | 47.7 / 47.7 / 47.7 | 52.5 / 50.8 / 51.3 |
+| Qwen3-VL-30B-A3B-Thinking | 62.7 / 61.0 / 61.0 | 65.1 / 65.1 / 65.1 | 65.2 / 65.2 / 65.2 | 75.4 / 74.6 / 73.7 | 78.9 / 78.4 / 78.9 | 17.1 / 17.1 / 17.1 | 75.8 / 75.8 / 75.8 | 61.5 / 61.5 / 61.5 | 65.0 / 64.7 / 64.6 |
+| Qwen3-VL-235B-A22B-Instruct | 76.3 / 72.5 / 74.6 | 68.9 / 67.2 / 68.0 | 63.6 / 63.6 / 63.6 | 80.7 / 79.4 / 80.7 | 75.0 / 74.0 / 75.0 | 8.6 / 8.6 / 8.6 | 74.2 / 74.2 / 74.2 | 56.9 / 56.5 / 56.9 | 66.0 / 64.9 / 65.6 |
+| Model | PD | AN | LC | AR | VC | CP | DM | FP | ALL |
+| AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | |
+| Closed-sourced MLLMs | |||||||||
+| GPT-5-Na no | 37.3 / - / - | 56.3 / - / - | 65.2 / - / - | 68.4 / - / - | 71.2 / - / - | 5.7 / - / - | 56.1 / - / - | 44.6 / - / - | 53.1 / - / - |
+| GPT-5 | 76.3 / - / - | 64.1 / - / - | 71.2 / - / - | 82.5 / - / - | 86.5 / - / - | 25.7 / - / - | 81.8 / - / - | 64.6 / - / - | 70.6 / - / - |
+| Gemini-2.5-Flash | 44.1 / 42.4 / 42.4 | 48.5 / 44.9 / 44.7 | 60.6 / 60.6 / 60.6 | 66.7 / 65.8 / 66.7 | 57.7 / 55.3 / 55.8 | 8.6 / 8.6 / 8.6 | 48.5 / 46.6 / 47.0 | 29.2 / 28.9 / 29.2 | 47.3 / 45.7 / 45.9 |
+| Gemini-2.5-Pro | 79.7 / 67.0 / 59.3 | 54.4 / 45.6 / 41.8 | 68.2 / 65.2 / 62.1 | 75.4 / 71.9 / 71.9 | 78.9 / 73.1 / 71.2 | 14.3 / 14.3 / 14.3 | 63.6 / 61.4 / 59.1 | 60.0 / 54.2 / 52.3 | 63.2 / 57.5 / 54.7 |
+| Claude-Sonnet-4.5 | 44.1 / 32.6 / 30.5 | 29.1 / 27.9 / 27.2 | 36.4 / 33.0 / 31.8 | 50.9 / 49.1 / 49.1 | 53.9 / 50.5 / 50.0 | 2.9 / 2.9 / 2.9 | 33.33 / 29.6 / 27.3 | 13.9 / 10.4 / 9.2 | 33.6 / 30.1 / 29.0 |
+| Open-sourced VLMs | |||||||||
+| LLaVA-OneVision-1.5-4B-Instruct | 6.7 / - / - | 4.9 / - / - | 10.6 / - / - | 12.3 / - / - | 5.8 / - / - | 0.0 / - / - | 7.6/ - / - | 0.0 / - / - | 6.2 / - / - |
+| LLaVA-OneVision-1.5-8B-Instruct | 15.3 / - / - | 1.9 / - / - | 7.6 / - / - | 1.8 / - / - | 1.9 / - / - | 0.0 / - / - | 4.6 / - / - | 3.1 / - / - | 4.6 / - / - |
+| GLM-4.5V | 54.2 / - / - | 28.2 / - / - | 30.4 / - / - | 59.7 / - / - | 51.9 / - / - | 5.7 / - / - | 54.6 / - / - | 6.2 / - / - | 37.4 / - / - |
+| GLM-4.1V-9B-Thinking | 35.6 / 32.6 / 32.2 | 18.5 / 18.0 / 17.5 | 33.3 / 28.4 / 27.3 | 42.1 / 39.9 / 42.1 | 38.5 / 35.1 / 36.5 | 0.0 / 0.0 / 0.0 | 30.3 / 28.0 / 28.8 | 7.7 / 6.2 / 4.6 | 26.0 / 23.9 / 23.9 |
+| Llama-3.2-90B-Vision-Instruct | 39.0 / 29.2 / 23.7 | 15.5 / 9.5 / 7.8 | 24.2 / 21.0 / 18.2 | 47.4 / 44.7 / 43.9 | 40.4 / 36.5 / 34.6 | 2.9 / 1.4 / 0.0 | 9.1 / 6.4 / 6.1 | 7.7 / 3.9 / 1.5 | 22.5 / 19.1 / 17.7 |
+| Llama-4-Maverick-17B-Instruct | 37.3 / 30.9 / 30.5 | 15.5 / 12.1 / 9.7 | 25.8 / 22.7 / 22.7 | 50.9 / 46.9 / 43.9 | 38.5 / 34.1 / 32.7 | 0.0 / 0.0 / 0.0 | 10.6 / 7.2 / 4.6 | 10.8 / 7.3 / 6.2 | 23.1 / 18.5 / 16.5 |
+| InternVL3-78B | 34.0 / 19.9 / 15.3 | 3.9 / 3.4 / 2.9 | 16.7 / 15.2 / 16.7 | 33.3 / 27.6 / 24.6 | 19.2 / 17.3 / 17.3 | 2.9 / 1.4 / 0.0 | 9.1 / 5.3 / 3.0 | 7.7 / 3.9 / 0.0 | 15.1 / 11.2 / 8.5 |
+| InternVL3.5-8B | 27.1 / 24.6 / 22.0 | 35.9 / 35.0 / 35.0 | 33.3 / 33.0 / 33.3 | 40.4 / 39.5/ 38.6 | 38.5 / 37.0 / 36.5 | 5.7 / 5.7 / 5.7 | 36.4 / 35.6 / 34.9 | 23.1 / 21.9 / 21.5 | 31.6 / 30.6 / 30.0 |
+| DeepSeek-VL2 | 11.9 / 5.9 / 0.0 | 1.0 / 0.1 / 0.1 | 7.6 / 4.2 / 1.5 | 10.5 / 8.3 / 7.0 | 3.9 / 3.4 / 3.9 | 2.9 / 1.4 / 0.0 | 3.0 / 1.5 / 0.0 | 6.2 / 2.3 / 0.0 | 5.6 / 3.0 / 1.6 |
+| Qwen3-VL-8B-Instruct | 33.9 / 31.4 / 32.2 | 47.6 / 46.4 / 46.6 | 37.9 / 36.7 / 37.9 | 50.9 / 48.3 / 49.1 | 38.5 / 38.5 / 38.5 | 2.9 / 2.9 / 2.9 | 54.6 / 53.4 / 53.0 | 29.2 / 27.7 / 29.2 | 39.6 / 38.2 / 38.8 |
+| Qwen3-VL-8B-Thinking | 59.3 / 59.3 / 59.3 | 60.2 / 60.2 / 60.2 | 60.6 / 60.6 / 60.6 | 64.9 / 64.5 / 64.9 | 63.5 / 63.5 / 63.5 | 11.4 / 10.0 / 8.6 | 74.2 / 73.9 / 74.2 | 47.7 / 47.3 / 47.7 | 57.9 / 57.6 / 57.7 |
+| Qwen3-VL-30B-A3B-Instruct | 35.6 / 32.6 / 33.9 | 59.2 / 58.5 / 59.2 | 51.5 /50.4/ 51.5 | 63.2 / 60.1 / 59.7 | 57.7 / 57.2 / 57.7 | 5.7 / 5.7 / 5.7 | 62.1 / 61.4 / 62.1 | 38.5 / 38.1 / 38.5 | 49.7 / 48.5 / 49.1 |
+| Qwen3-VL-30B-A3B-Thinking | 74.6 / 74.6 / 74.6 | 65.1 / 65.1 / 65.1 | 66.7 / 66.7 / 66.7 | 71.9 / 71.9 / 71.9 | 78.9 / 78.4 / 78.9 | 5.7 / 5.7 / 5.7 | 83.3 / 83.0 / 83.3 | 66.2 / 66.2 / 66.2 | 67.0 / 66.9 / 67.0 |
+| Qwen3-VL-235B-A22B-Instruct | 72.9 / 71.2 / 72.9 | 58.3 / 57.5 / 57.3 | 59.1 / 58.3 / 59.1 | 73.7 / 73.7 / 73.7 | 71.2 / 69.7 / 69.2 | 5.7 / 5.7 / 5.7 | 60.6 / 59.1 / 59.1 | 58.5 / 58.1 / 58.5 | 59.8 / 59.0 / 59.2 |
+Appendix D More Details of Experiment Results
+In this section, we provide more details of the experiment results.
+D.1 More Details of Experimental Results on Text-Image Modality
+In Section 4.2, we presented the evaluation results for three batches of instances sampled with random seeds 0, 1, and 2.
+Here, we further present the performance of various VLMs on each batch of sampled instances, as shown in Tables 6, 7, and 8.
+It can be observed that the closed-source models exhibit minimal fluctuations across the three sampled variant instances, with changes in Answer Accuracy (AA) remaining under 4%, demonstrating strong robustness.
+Among the open-source models, all except for the Qwen3-VL family show similar stability, likely because the varying difficulty of the tasks has little impact on these weaker models, as they are unable to solve the problems in the first place.
+In contrast, for the Qwen3-VL family, except for Qwen3-VL-30B-A3B-Instruct and Qwen3-VL-30B-A3B-Thinking, other models show greater fluctuations in performance.
+Specifically, Qwen3-VL-8B-Instruct shows a 6.7% change in AA, Qwen3-VL-8B-Thinking shows a 9.3% change, and Qwen3-VL-235B-A22B-Instruct shows a 6.2% change.
+This suggests that while the Qwen3-VL family models generally perform on par with some closed-source models, their robustness still requires improvement.
+| Model | PD | AN | LC | AR | VC | CP | DM | FP | ALL |
+| AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | AA / PS / PA | |
+| Qwen3-VL-8B-Instruct | 16.9 / 16.9 / 16.9 | 35.9 / 35.2 / 35.9 | 28.8 / 27.7 / 28.8 | 24.6 / 24.1 / 24.6 | 34.6 / 34.1 / 34.6 | 0.0 / 0.0 / 0.0 | 43.9 / 43.2 / 43.9 | 16.9 / 16.9 / 16.9 | 27.4 / 26.9 / 27.4 |
+| Qwen3-VL-8B-Thinking | 35.6 / 35.6 / 35.6 | 44.7 / 44.7 / 44.7 | 36.4 / 36.4 / 36.4 | 47.4 / 47.4 / 47.4 | 28.9 / 28.9 / 28.9 | 2.9 / 2.9 / 2.9 | 45.5 / 45.5 / 45.5 | 18.5 / 18.5 / 18.5 | 35.0 / 35.0 / 35.0 |
+| Qwen3-VL-30B-A3B-Instruct | 32.2 / 29.2 / 30.5 | 40.8 / 40.5 / 40.8 | 40.9 / 40.5 / 40.9 | / 61.4 / 61.4 | 40.4 / 39.9 / 40.4 | / 8.6 / 8.6 | 59.1 / 59.1 / 59.1 | 24.6 / 24.2 / 24.6 | 40.2 / 39.6 / 40.0 |
+| Qwen3-VL-30B-A3B-Thinking | / 54.2 / 54.2 | / 53.4 / 53.4 | / 47.0 / 47.0 | 47.4 / 47.4 / 47.4 | / 48.1 / 48.1 | / 8.6 / 8.6 | / 62.1 / 62.1 | / 36.9 / 36.9 | / 47.3 / 47.3 |
+D.2 Experimental Results on Text-Video Modality
+Here, we further evaluate VLMs on the text-video modality of DynaSolidGeo.
+Since the OpenRouter API does not yet support video inputs for most MLLMs–and even for the Gemini family, the support remains unstable (frequent network errors occur during API calls)–we therefore locally deploy and evaluate several representative open-source models: Qwen3-VL-8B-Instruct, Qwen3-VL-8B-Thinking, Qwen3-VL-30B-A3B-Instruct, and Qwen3-VL-30B-A3B-Thinking.
+In particular, we conduct the evaluation on a batch of instances sampled with random seed 0, while keeping all model hyperparameters consistent with those in Section 4.1.
+The results are summarized in Table 9.
+Overall, these VLMs perform worse on the text-video modality compared to the text-image modality.
+Across different task categories, the best performance on Area calculation (AR) is achieved by Qwen3-VL-30B-A3B-Instruct, with AA, PS, and PA all reaching 61.4%.
+In all other categories, Qwen3-VL-30B-A3B-Thinking performs best.
+On Counting Problems (CP), all models perform extremely poorly, with all metrics below 9%, and Qwen3-VL-8B-Instruct even drops to 0.0%, revealing a major limitation of current VLMs on this task.
+Furthermore, the thinking models show smaller drops in PA relative to AA compared to instruct models, indicating stronger reasoning capability, which is consistent with the trend observed in the text-image modality.
+D.3 More Details of Error Analysis
+Here, we provide additional details of the error analysis conducted on two representative models: Gemini-2.5-Pro and Qwen3-VL-30B-A3B-Thinking.
+Specifically, we thoroughly examine the distribution of error types–Visual Perception Errors, Logical Reasoning Errors, Calculation Errors, Knowledge Errors, and Hallucination Errors–across different task categories and difficulty levels.
+The results are presented in Figures 5 and 6.
+For Gemini-2.5-Pro, Visual Perception Errors and Logical Reasoning Errors dominate across most categories.
+In Positional Relationship Determination (PD), Angle Calculation (AN), Length and Distance Calculation (LC), Volume Calculation (VC), Counting Problems (CP), and Folding and Unfolding Problems (FP), the proportion of Visual Perception Errors for Gemini-2.5-Pro gradually decreases as the difficulty level increases, while the share of Logical Reasoning Errors rises accordingly.
+This suggests that when reasoning cues become more complex, Gemini-2.5-Pro maintains relatively stable perception but struggles to reason through multi-step or abstract geometric relations.
+Moreover, Hallucination Errors start to emerge in medium- and high-difficulty problems, implying that the model tends to compensate for uncertainty in perception or reasoning by fabricating unsupported intermediate steps, revealing its limited robustness in handling complex 3D spatial reasoning.
+By contrast, Qwen3-VL-30B-A3B-Thinking exhibits a much higher proportion of Visual Perception Errors but markedly fewer reasoning and hallucination errors, indicating that it maintains more consistent logical structures once visual understanding is correct.
+This finds that Qwen3-VL-30B-A3B-Thinking is stronger in symbolic reasoning but more sensitive to visual ambiguity or occlusion.
+Overall, the two models display complementary weaknesses: Gemini-2.5-Pro tends to misreason over correct visual cues, while Qwen3-VL-30B-A3B-Thinking tends to misperceive geometric relations even when reasoning remains sound.
+These trends further confirm that current VLMs still struggle to integrate visual perception and logical reasoning coherently in 3D spatial problem solving.
+Appendix E DynaSolidGeo as a Training Dataset
+To assess the effectiveness and practical utility of DynaSolidGeo, we additionally examine its impact on VLM performance when used as a training dataset.
+E.1 Experimental Setup
+We sample batches of instances using random seeds from 0 to 9, resulting in a total of 5,030 samples.
+These samples are divided into a training set (3,627 samples), a validation set (403 samples), and a test set (1,000 samples).
+The distribution across task categories is kept consistent among the three subsets, and all subsets are ensured to be non-overlapping with respect to the seed questions.
+Using the training set, we fine-tuned Qwen3-VL-4B-Instruct and Qwen3-VL-8B-Instruct with the Group Relative Policy Optimization (GRPO) algorithm, obtaining Qwen3-VL-4B-DynaSolidGeo and Qwen3-VL-8B-DynaSolidGeo, respectively.
+We then compared the changes in their Answer Accuracy (AA) before and after training.
+We conduct GRPO training using the Verl framework on 8*NVIDIA H100 GPUs.
+Specifically, we adopt an answer-based reward scheme, assigning a reward of 1 for correct answers and 0 otherwise.
+Each question is rolled out 5 responses, with a maximum response length of 4096 tokens.
+For other hyperparameters, we set the learning rate to 1e-6, the batch size to 512, and train for 3 epochs, totaling 21 steps.
+E.2 Experimental Results
+The experimental results are shown in Table 10.
+After being fine-tuned on the DynaSolid training set, the performances of Qwen3-VL-4B-Instruct and Qwen3-VL-8B-Instruct improved by 8.6% and 9.0%, respectively.
+The remarkable performance gain highlights the effectiveness and high quality of DynaSolidGeo dataset.
+As a large-scale dynamic benchmark, it establishes a new foundation for evaluating and improving spatial mathematical reasoning in multimodal models, offering the community a robust and extensible resource for advancing reasoning-centric VLM research.
+| Model | Answer Accuracy (AA) |
+| Qwen3-VL-4B-Instruct | 16.2 |
+| Qwen3-VL-4B-DynaSolidGeo | 24.8 (+8.6) |
+| Qwen3-VL-8B-Instruct | 17.0 |
+| Qwen3-VL-8B-DynaSolidGeo | 26.0 (+9.0) |
+Appendix F More Details of Data Annotation Pipeline
+In this section, we provide more details of data annotation pipeline.
+F.1 Parameterized Question Example
+In the JSON-Based Question Parametrization stage, the variable parameters in the question statements (e.g., endpoints, side lengths, areas, volumes, ratios) are represented as variables enclosed in curly braces {} using f-string syntax.
+For example, the endpoint “A” is replaced with {point_A}, and the side length is replaced with {len_a}.
+In this way, during the subsequent Parameterized Program Construction stage, these placeholders can be directly assigned to Python variables when assembling the parameterized program.
+An example of the parameterized question annotation (in JSON format) is shown below.
+F.2 Answer Function Example
+To ensure the dynamic generation of ground-truth answers according to the variable parameters defined in each question statement, a dedicated Python function (Answer Function) is created for every seed question.
+Each function computes the ground truth using the annotated variables from the parameterized question.
+An example of the annotated answer function is provided below.
+F.3 Parameterized Python Program Example
+In the Parameterized Program Construction stage, we assemble the Parameterized Python Program by combining the parameterized question with its corresponding answer function.
+An example is shown below.
+The program randomizes question variables and image camera viewpoints.
+F.4 Parameterized Visualization Program Example
+In the MATLAB-Based Geometry Visualization stage, the parameterized question and the static MATLAB visualization program are integrated into a Parameterized Visualization Program.
+An example is shown below.
+It takes the variable parameters (including camera view angles) passed from the Parameterized Python Program and dynamically renders the image or video.
+F.5 Prompts for Data Annotation Pipeline
+In the Expert-Guided Parametrization and Visualization stage of the data annotation pipeline, we follow a human-in-the-loop strategy, where human experts collaborate with large models to create a JSON annotation and a MATLAB program for each source question.
+The prompts for JSON generation and MATLAB code generation are shown below.
+Appendix G Question Instance Examples
+In this section, we provide several examples of question instances, as shown in the text boxes at the end of this appendix.
+For each task category, we present three variations sampled with random seeds 0, 1, and 2, with each variable parameter highlighted in red.